You are a farmer and want to determine whether sheep wool production (measured in fleece weight) is a heritable trait. A survey of 800 sheep from the surrounding area found the average fleece weight to be 7.2 lbs.

You selected a group of the 40 most productive sheep (average fleece weight 9.3 lbs), which are later bred, and the wool production of their offspring (at maturity) is measured. You found the average fleece weight of the offspring to be 8.19 lbs.

a. Calculate the selection differential

b. Calculate the response to selection.

c. Calculate heritability (h2).  

-If you were to carry out this breeding program for multiple generations would you expect to see wool production to be strongly influenced by genetics, environment, or a combination?
Answer #1

a) Selection differential = Mean of selected parents from base population - Mean of base population

= 9.3 - 7.2

= 2.1

b) response to selection = Mean of offspring of the selected parents - Mean of base population

= 8.19 - 7.2

= 0.99

c) Heritability (h​​​​​​2​​​​​​​​​​) = Selection response ÷ Selection differential

= 0.99÷2.1

= 0.47

Hence about half of the phenotype of the offspring is attributable to solely genes. The rest is due to other factors including the environment. Hence wool production will be strongly influenced by a combination of both genetics and environment.